OpenStudy (anonymous):

Can someone check these matrices, please. Solve each matrix equation 7.[2 -6 8] + [-1 -2 4] = X row reduce. [1 0 -4/5] [0 1 -8/5] 9. [ x– 5 9] = [-7 w+1] [2] [8] [4 t+2] [8-r 1] [4] [-1]

OpenStudy (anonymous):

5.[ 1 2 -5] + [-2 7 -3] = [-2+1 7+9 -5+-3] = [-1 9 -8] [3 -2 1] [1 2 5] [3+1 -2+2 1+5] = [4 0 6]

OpenStudy (anonymous):

[x - 5]